Greg owns and operates a florist. Greg faces the linear inverse demand function P(Q) = 37 – 0.5Q. If follows that Greg faces the marginal revenue function,a)MR(Q) = 37 – 0.25Q.b)MR(Q) = 74 – 0.5Q.c)MR(Q) = 37 - Q.d)MR(Q) = 37 - 2Q.
Answer #1

Marginal revenue = d(TR) / dQ

TR = P*Q = (37 - 0.5Q)Q = 37Q - 0.5Q2

MR = d( 37Q - 0.5Q2) / dQ = 37 - 0.5*2Q

MR = 37 - Q

So, Greg faces the marginal revenue function MR (Q) = 37 - Q

correct ans is c) MR(Q) = 37 - Q